Find all the 5th roots of (5-i).
1
Expert's answer
2020-12-16T20:10:08-0500

"5-i=\sqrt{26}e^{i\phi}" , "\phi=\sin^{-1}({-\frac{1}{\sqrt{26}}})"

if "z^5=5-i" then

"z=26^{\frac{1}{10}}e^{\frac{i\phi}{5}+\frac{2k\pi}{5}} , k=0, 1, 2, 3, 4".